# Disentangling Representations in RNNs through Multi-task Learning

### v1

@@ Abstract
Abstract, or disentangled, representations are a promising mathematical framework for efficient and effective generalization in both biological and artificial systems. We investigate abstract representations in the context of multi-task classification over noisy evidence streams – a canonical decision-making neuroscience paradigm. We derive theoretical bounds that guarantee the emergence of disentangled representations in the latent state of any optimal multi-task classifier, when the number of tasks exceeds the dimensionality of the state space. We experimentally confirm that RNNs trained on multi-task classification learn disentangled representations in the form of continuous attractors, leading to zero-shot out-of-distribution (OOD) generalization. We demonstrate the flexibility of the abstract RNN representations across various decision boundary geometries and in tasks requiring classification confidence estimation. Our framework suggests a general principle for the formation of cognitive maps that organize knowledge to enable flexible generalization in biological and artificial systems alike, and closely relates to representations found in humans and animals during decision-making and spatial reasoning tasks.
@@ 1 Introduction
Humans and animals can generalize to new settings effortlessly, leveraging a combination of past experiences and world models [43, 44]. Modern foundation models also display emergent out-of-distribution (OOD) generalization abilities, in the form of zero- or few-shot learning [10, 59, 36, 57]. For generalization, it is crucial to have appropriate representations which extrapolate or interpolate previous experiences [67, 35, 38], or the use of task compositionality as an inductive bias [76, 62].
One mechanism for generalization is through abstract, or disentangled, representations [29, 39, 38]. A property of abstract representations is that they preserve the latent structure present in the world in their geometry, i.e. they are low-dimensional, rectangular, and compositional, each axis corresponding to a factor of variation in the data. Then a linear decoder (i.e. downstream neuron) trained to discriminate between two categories can readily generalize to stimuli not observed in training, due to the structure of the representation. From a neuroscience perspective, this corresponds to decomposing a novel stimulus into its familiar attributes, and reasoning based on those. For instance, imagine you are at a grocery store, deciding whether a fruit is ripe or not. If the brain’s internal representation of food attributes (ripeness, caloric content, etc.) is disentangled, then learning to perform this task for bananas would lead to zero-shot generalization to other fruit (e.g. mangos, Figure 1a).
This raises the question of under which conditions do such representations emerge in biological and artificial agents alike. Several brain areas including the amygdala, prefrontal cortex and hippocampus have been found to encode variables of interest in an abstract format [65, 5, 8, 55, 16]. Previous work showed that feedforward neural networks develop abstract representations when trained to multitask [38]. However, real-world decisions typically rely on imperfect, noisy information, evolving dynamically over time [9, 41]. To account for this important feature of the world, we train RNNs to multitask canonical neuroscience tasks involving the accumulation of evidence over noisy streams.
@@
The main contributions of this paper are the following:
  • We prove that any optimal multi-task classifier is guaranteed to learn a disentangled representation of the ground truth contained in the noisy measurements in its latent state, if the classification boundary normal vectors span the input space (Appendix B). Intriguingly, noise in the observations is necessary to guarantee the latent state would compute an optimal, disentangled representation of the ground truth.
  • Through experiments, we confirm that RNNs trained to multitask develop abstract representations that zero-shot generalize OOD, when the number of tasks exceeds the input dimensionality. The computational substrate of these representations is a 2D continuous attractor [2] storing a ground truth estimate in a product space of the latent factors.
  • We demonstrate that our setting is robust to a number of manipulations, including interleaved learning of linear and non-linear tasks and free reaction time decisions.
  • Finally, we discuss the relation between our results and important neuroscientific findings, their implications for generalizable representation learning in artificial systems, and demonstrate the strong advantage of multitasking over previously proposed mechanisms of representation learning in the brain [48, 76].
Despite being framed in the context of neuroscience, our results are general; they apply to any system aggregating noisy evidence over time.
@@ 1.1 Related work
Disentanglement has long been recognized as a promising strategy for generalization [4] (although note [46, 54] for a contrarian view), yet most modern work focuses on feedforward architectures [29, 39, 74]. In terms of autoregressive models, Hsu et al. [33], Li and Mandt [45] showed that variational LSTMs disentagle representations of underlying factors in sequential data allowing style transfer; however to the best of our knowledge the underlying representational geometry has not been characterised. Other work focuses on fitting RNNs to behavioral data while enforcing disentanglement for interpretability [18, 51]. Work on context-dependent decision making has shown that RNNs re-purpose learned representations in a compositional manner when trained in related tasks [76, 20]; however, the abstractness of the resulting representations was not established. Finally John et al. [37] show that multitasking results in disentanglement, however they directly enforce latent factor separation through their adversarial optimization objectives. Our approach is most closely related to weakly supervised disentanglement, without comparing across samples [66].
@@
Previous work showed that multitasking feedforward networks learn abstract representations, as quantified by regression generalization [38]. We expand upon these findings in several ways. First, we extend the framework of Johnston and Fusi [38] to RNNs that can update their representations as further information arrives. Second, we prove theorems on optimal multi-task classifiers that guarantee the emergence of disentangled representations in any optimal multitask classifier if the number of tasks exceeds the input dimensionality $D$. Third, we rigorously analyze the role of noise in forming disentangled representations, extending the noise-free regime studied in [38]. Finally, we explore a range of values for $D$, providing experimental validation of our theory. Whereas related work proved that disentangled representations emerge in feedforward architectures from multitask learning in sparse tasks when a sparsity regularization constraint is placed on the predictors [42], we place no such constraints and still uncover disentangled representations.
2 Model and setup
2.1 Problem formulation
We consider canonical cognitive neuroscience tasks that involve evidence aggregation over time, mirroring decision-making under uncertainty. The input consists of $D$ noisy evidence streams $\mathbf{X}(t)\in\mathbb{R}^{D}$ where $\mathbf{X}(t)=\mathbf{x}^{*}+\sigma\,\mathcal{N}(\mathbf{0},\mathbf{I}_{D})$ and $\mathbf{x}^{*}\in\mathbb{R}^{D}$ is the ground truth evidence in a certain trial ($x_{i}^{*}\sim\textit{Uniform}(-0.5,0.5)$) and $\sigma$ is the input noise standard deviation (Figure 1c, left). Each element $x_{i}^{*}$ of $\mathbf{x}^{*}$ corresponds to different options under consideration or different attributes of the same item. The target output $\mathbf{y}(\mathbf{x}^{*})\in\{-1,+1\}^{N_{\text{task}}}$ is a vector of $N_{\text{task}}$ +1s and -1s, depending on whether $\mathbf{x}^{*}$ is above or below each of the classification boundaries (Figure 1b). To ensure the reproducibility of our results, we incorporate our tasks within the NeuroGym framework [53].
@@ 2.1 Problem formulation
The decision maker should integrate noisy samples $\mathbf{X}(t)$ over time, viewed through an injective observation map (encoder) $f$, to estimate $\hat{\mathbf{Y}}_{i}(t)=\Pr\{\mathbf{y}_{i}(\mathbf{x}^{*})=1|f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))\}$ (Figure S2). The classification lines reflect criteria based on which decisions will be made. Imagine for example that $x_{1}$ corresponds to food and $x_{2}$ to water reward. Depending on the agent’s internal state, one could take precedence over the other, and the degree of preference is reflected in the slope of the line. In Appendix B we prove that any optimal multi-task classifier will learn disentangled representations of $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$, when the $N_{\text{task}}$ classification boundaries span the latent space.
2.2 Architecture
We train leaky RNNs which represent a brain area responsible for decision-making to multitask the aforementioned tasks. The networks contain $N_{\textit{neu}}=64$ neurons, and their activations $\mathbf{z}(t)$ obey:
$\tau\frac{{\textrm{d}}\mathbf{z}(t)}{{\textrm{d}}t}=-\mathbf{z}(t)+\left[\,\mathbf{W}_{\textit{rec}}\,\mathbf{z}(t)+\mathbf{W}_{\textit{in}}\,\mathbf{x}_{\textit{in}}(t)+\mathbf{b}\,\right]_{+}$ (1)
where $\mathbf{W}_{\textit{rec}}$ is the recurrent weight matrix, $\mathbf{W}_{\textit{in}}$ is the matrix carrying the input vector $\mathbf{x}_{\textit{in}}$, $\mathbf{b}$ is a unit-specific bias vector, $\tau$ is the neuronal time constant and $\left[.\right]_{+}$ is the ReLU applied element-wise to a vector. We simulate Equation 1 using the forward Euler method for $T=20$ timesteps of duration $\Delta t=\tau=100\,\textrm{ms}$, which we find to be stable. The RNN’s output $\hat{\mathbf{y}}(t)$ of size $N_{\textit{task}}$ is given by:
$\hat{\mathbf{y}}(t)=g(\mathbf{W}_{\textit{out}}\,\mathbf{z}(t))$ (2)
where $\mathbf{W}_{\textit{out}}$ is a readout matrix and $g$ the output activation function applied elementwise. We choose $g(z)=tanh(z)$ for classification, $5\,tanh(z)$ for integration-to-bound, and $g(z)=z$ for other tasks. The inputs described above are first passed through a randomly initialized encoder $f$, which represents a static mapping from latents to observations (e.g. a perceptual system). The encoder is a 3-layer MLP with hidden dimensions $100,100,40$ and ReLU non-linearities, is unique to each network and kept fixed during training. An additional fixation input which is 1 during the trial and turns 0 when the network should report its decisions is directly passed to the hidden layer (Figure 1c). The RNN is trained to produce the target outputs. By minimizing loss across trials, the network is incentivized to estimate $\Pr\{y_{i}(\mathbf{x}^{*})=+1\}$. We use mean-squared-error loss and train all parameters (i.e. $\mathbf{W}_{\textit{rec}}$, $\mathbf{W}_{\textit{in}}$, $\mathbf{W}_{\textit{out}}$ and $\mathbf{b}$) with Adam default settings except $\text{lr}=10^{-3}$ [40]. Table S1 summarizes all hyperparameters and their values. Network training takes $\sim 10$ minutes on a commercial laptop CPU.
@@ 3 Theoretical Results
We seek to understand the properties of optimal multi-task classifiers with random variable latent state $\mathbf{Z}(t)$ in the paradigm illustrated in Figure S2 and described in Section 2.1. Specifically, we are interested in classifiers that, at each time point $t$, estimate multiple linear classifications $\mathbf{y}(\mathbf{x}^{*})=[y_{1}(\mathbf{x}^{*}),\dots,y_{N_{\rm task}}(\mathbf{x}^{*})]\in\{0,1\}^{N_{\rm task}}$ based on noisy, potentially non-linearly mapped observations $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$ with maximal likelihood. The classifier output at time $t$, denoted $\hat{\mathbf{Y}}(t)\in[0,1]^{N_{\rm task}}$, represents a set of Bernoulli random variable parameters, where each element $\hat{Y}_{i}(t)$ corresponds to the probability $\Pr\{y_{i}(\mathbf{x}^{*})=1|f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))\}$).
Each of the $N_{\text{task}}$ decision boundaries can be expressed as a pair of boundary normal vector $\mathbf{c}_{i}\in\mathbb{R}^{D}$ and offset $b_{i}\in\mathbb{R}$ (Equation 8). Let $\mathbf{C}\in\mathbb{R}^{N_{task}\times D}$ represent the matrix of decision boundaries and $\mathbf{b}\in\mathbb{R}^{N_{task}}$ be the vector of decision boundary offsets.
@@
If $\mathbf{C}\in\mathbb{R}^{N_{task}\times D}$ is a full-rank matrix and $N_{task}\geq D$, then
@@
    any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode an estimate $\mu(t)$ of the ground truth evidence variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$, and
@@
    if the activation function $g=\tanh$, $\mu(t)$ will be linearly decodable from $\mathbf{Z}(t)$, thus implying that $\mathbf{Z}(t)$ contains a linear disentangled representation of $\mu(t)$ [30, 38].
Specifically, $\mu(t)$ is the maximum likelihood estimate of $\mathbf{x}^{*}$ given observations $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$. Equations 34 are closed-form expressions for extracting $\mu(t)$ from $\mathbf{Z}(t)$ with a generic activation function $g$ and $g=\tanh$ respectively.
@@
$\approx\underbrace{\frac{2\sqrt{3}\sigma}{\pi\sqrt{t}}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{Z}(t)}_{\text{Linear Function of $\mathbf{Z}(t)$}}+\underbrace{(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{b}}_{\text{Affine Term}}$ (4)
Where $\Phi$ is the CDF of the normal distribution, $\sigma$ is the noise magnitude and $t$ the trial duration.
@@
Point 1 and Equation 3 are proven in Appendix B in Theorem B.6. Point 2 and Equation 4 are proven in Corollary B.7. ∎
Theorem 3.1 holds for any system that performs optimal multi-task classification based on noisy, injectively transformed observations $f(\mathbf{X}(t))$ (e.g., RNNs, biological neural circuits, Bayesian filters, etc.). Our results are readily extensible to sub-optimal multi-task classification models under the assumption that classification error w.r.t. optimal estimation is randomly distributed with zero mean (see Appendix B). In this case, the closed-form solutions in Equations 34 correspond to a least-squared error estimate of $\mu(t)$ given sub-optimal $Z(t)$. Surprisingly, our theoretical results indicate $\mathbf{Z}(t)$ will contain a disentangled representation of $\mathbf{x}^{*}$ when the readout map $g(\mathbf{Z}(t)):=\hat{\mathbf{Y}}(t)$ is a $\tanh$ function (Equation 4, Corollary B.7). Our theoretical framework readily generates testable predictions for our computational experiments, including theoretical $r^{2}$ values for estimating $\mathbf{x}^{*}$ from $\mathbf{Z}(t)$ (Appendix A.3, Figure S9a), and decodability of $\mathbf{x}^{*}$ from $\mathbf{Z}(t)$ as a function of latent dimension $D$ (Figure 3a), number of tasks $N_{task}$ (Figure 1f,g, 2c), noise level $\sigma$ (Figure 3b), evidence aggregation time $t$ (Lemma B.3), and latent factor correlation (Figure S6).

### v2

# Disentangling Representations through Multi-task Learning

@@ Abstract
Intelligent perception and interaction with the world hinges on internal representations that capture its underlying structure (“disentangled” or “abstract” representations). Disentangled representations serve as world models, isolating latent factors of variation in the world along approximately orthogonal directions, thus facilitating feature-based generalization. We provide experimental and theoretical results guaranteeing the emergence of disentangled representations in agents that optimally solve multi-task evidence aggregation classification tasks, canonical in the cognitive neuroscience literature. The key conceptual finding is that, by producing accurate multi-task classification estimates, a system implicitly represents a set of coordinates specifying a disentangled representation of the underlying latent state of the data it receives. The theory provides conditions for the emergence of these representations in terms of noise, number of tasks, and evidence aggregation time. Surprisingly, the theory also produces closed-form expressions for extracting the disentangled representation from the model’s latent state $\mathbf{Z}(t)$. We experimentally validate these predictions in RNNs trained on multi-task classification, which learn disentangled representations in the form of continuous attractors, leading to zero-shot out-of-distribution (OOD) generalization in predicting latent factors. We demonstrate the robustness of our framework across autoregressive architectures, decision boundary geometries and in tasks requiring classification confidence estimation. We find that transformers are particularly suited for disentangling representations, which might explain their unique world understanding abilities. Overall, our framework puts forth parallel processing as a general principle for the formation of cognitive maps that capture the structure of the world in both biological and artificial systems, and helps explain why ANNs often arrive at human-interpretable concepts, and how they both may acquire exceptional zero-shot generalization capabilities.
@@ 1 Introduction
The ability to construct representations that capture the underlying structure of the world from data, is a hallmark of intelligence. Humans and animals leverage their experiences to construct such faithful representations of the world ("world models"), resulting in a near-effortless ability to generalize to new settings [42, 43]. Modern foundation models also display emergent out-of-distribution (OOD) generalization abilities, in the form of zero- or few-shot learning [10, 61, 35, 58]; however whether artificial systems learn world models remains unclear. Understanding the conditions under which that occurs is bound to lead to better generalizable systems, and explain why artificial systems often converge to human interpretable, aligned representations of the world [71].
A promising direction towards understanding the construction of world models is abstract, or disentangled representations [30, 38, 37]. These two concepts are interrelated yet somewhat distinct (see definitions adapted from Ostojic and Fusi [59] in Section A.1). Shortly, an abstract representation of $x_{1},\dots,x_{n}$ represents each $x_{i}$ linearly and approximately mutually orthogonally. Disentangled representations encode each $x_{i}$ orthogonally, without the necessity of linearity. Both representations preserve the latent structure present in the world in their geometry by isolating factors of variation in the data, which facilitates downstream generalization. When a representation is abstract, a linear decoder (i.e. downstream neuron) trained to discriminate between two categories can readily generalize to stimuli not observed in training, due to the structure of the representation. Furthermore, the more disentangled the representation is, the lower the interference from other variables and hence the better the performance. This corresponds to decomposing a novel stimulus into its familiar features, and performing feature-based generalization. For instance, imagine you are at a grocery store, deciding whether a fruit is ripe or not. If the brain’s internal representation of food attributes (ripeness, caloric content, etc.) is disentangled, then learning to perform this task for bananas would lead to zero-shot generalization to other fruit (e.g. mangos, Figure 1a). Crucially, the visual representation of a mango is high-dimensional, non-linear and noisy, making it particularly challenging to extract a low dimensional latent like "ripeness".
Several brain areas including the amygdala, prefrontal cortex and hippocampus encode variables of interest in an abstract format [67, 5, 8, 56, 16]. This raises the question of under which conditions do such representations emerge in biological and artificial agents alike. Previous work showed that feedforward neural networks develop abstract representations when trained to multitask [37]. However, real-world decisions typically rely on imperfect, noisy information, evolving dynamically over time [9, 40]. To account for this important feature of the world, we train autoregressive models (RNNs, LSTMs, transformers) to multitask canonical neuroscience tasks involving the accumulation of evidence over noisy streams. The tasks tie closely to Bayesian filtering theory, and should be solved by any agent that deals with a noisy world.
@@
The main contributions of the paper are the following:
  • We prove that any optimal multi-task classifier is guaranteed to learn an abstract representation of the ground truth contained in the noisy measurements in its latent state, if the classification boundary normal vectors span the input space (Appendix B). Furthermore, the representations are guaranteed to be disentangled as the number of tasks $N_{\text{task}}$ greatly exceeds the input dimensionality $D$. Intriguingly, noise in the observations is necessary to guarantee the latent state would compute a disentangled representation of the ground truth.
  • We confirm that RNNs trained to multitask develop abstract representations that zero-shot generalize OOD, when $N_{\text{task}}\geq D$, and orthogonal, disentangled representations for greater $N_{\text{task}}$. The computational substrate of these representations is a 2D continuous attractor [3] storing a ground truth estimate in a product space of the latent factors. In addition, the representations are sparse and mixed, attributes of biological neural networks.
  • We reproduce these findings in GPT-2 models, which generalize better due to them learning disentangled representations already from $N_{\text{task}}\geq D$, confirming their appropriateness for constructing disentangled world models.
  • We demonstrate that our setting is robust to a number of manipulations, including correlated inputs, interleaved learning of tasks and free reaction-time tasks canonical in the cognitive neuroscience literature [9, 40].
  • Finally, we discuss implications for generalizable representation learning in biological and artificial systems, and demonstrate the strong advantage of multi-task learning over previously proposed mechanisms of representation learning in the brain [49].
Despite being framed in the context of canonical cognitive neuroscience tasks, our results are general; they apply to any system aggregating noisy evidence over time.
@@ 1.1 Related work
Disentanglement has long been recognized as a promising strategy for generalization [4] (although note Locatello et al. [47], Montero et al. [55] for a contrarian view), yet most classic work focuses on feedforward architectures [30, 38, 76]. In terms of autoregressive models, Hsu et al. [33], Li and Mandt [46] showed that variational LSTMs disentagle representations of underlying factors in sequential data allowing style transfer; however the underlying representational geometry was not characterised. Other work focuses on fitting RNNs to behavioral data while enforcing disentanglement for interpretability [18, 53]. Work on context-dependent decision making has shown that RNNs re-purpose learned representations in a compositional manner when trained in related tasks [78, 20]; however, the abstractness of the resulting representations was not established. Finally John et al. [36] show that multitasking results in disentanglement, however unlike us they directly enforce latent factor separation through their adversarial optimization objectives. Our approach is most closely related to weakly supervised disentanglement, without comparing across samples [68].
Our work relates to previous work on linear identifiability. Roeder et al. [64] show that representations of models trained on the same distribution must be linear transformations of each other; yet we go beyond their results to show that abstract representations are guaranteed to emerge under moderate conditions, irrespectively of the dimensionality of the input and model architecture. Lachapelle et al. [41] proved that disentangled representations emerge in feedforward architectures from multitask learning in sparse tasks when a sparsity regularization constraint is placed on the predictors; we place no such constraints and still uncover disentangled representations.
@@
Previous neuroscience-inspired work showed that multitasking feedforward networks learn abstract representations, as quantified by regression generalization [37]. We expand upon these findings in several ways. First, we extend the framework to autoregressive architectures (RNNs, LSTMs, transformers) that can update their representations as further information arrives. Second, we prove theorems that guarantee the emergence of abstract representations in any optimal multitask classifier if the number of tasks exceeds the input dimensionality $D$, and showcase disentanglement in our trained networks. Third, we rigorously analyze the role of noise in forming disentangled representations, extending the noise-free regime studied in Johnston and Fusi [37]. Finally, we explore a range of values for $D$, providing experimental validation of our theory.
2 Problem formulation
Multi-Task Classification with Evidence-Aggregation:
We study the evidence aggregation multi-task classification paradigm shown in Figure 1b. An agent with latent state $\mathbf{Z}(t)$ receives noisy, non-linearly mapped observations $\{f\big{(}\mathbf{X}(t)\big{)}\}_{t=1}^{T}$ where each $\mathbf{X}(t)=\mathbf{x}^{*}+\sigma\,\mathcal{N}(0,I_{D})$ is a noisy measurement of unknown ground truth vector $\mathbf{x}^{*}\in\mathbb{R}^{D}(x_{i}^{*}\sim\textit{Uniform}(-0.5,0.5))$, with $\mathcal{N}$ being Gaussian noise. The noisy measurements are transformed by an injective observation map function $f$, which can be non-linear and high dimensional, representing the wide range of sensory transformations found in real-world scenarios. The agent is tasked with simultaneously solving $N_{\textit{task}}$ classification problems by aggregating information over time, each defined by a random linear decision boundary [^1] Due to the observation map $f$, the tasks may appear non-linear from the perspective of the multi-task classification agent. in the ground truth space $\mathbb{R}^{D}$ i.e.
$y_{i}(\mathbf{x}^{*})=\begin{cases}1&\text{ if }\mathbf{c}_{i}^{\top}\mathbf{x}^{*}>b_{i}\\ 0&\text{ otherwise }\end{cases}$ (1)
@@ 2 Problem formulation
where $\mathbf{y}(\mathbf{x}^{*})\in\{0,1\}^{N_{\textit{task}}}$ represents the $N_{\textit{task}}$ classifications of $\mathbf{x}^{*}$, $\{(\mathbf{c}_{i},b_{i})\}_{i=1}^{N_{\textit{task}}}$ are the classification boundary normal vectors and offsets, and let $\hat{\mathbf{Y}}(t)=g(\mathbf{Z}(t))\in[0,1]^{N_{\textit{task}}}$ represent the agent’s predicted likelihood of $y_{i}(\mathbf{x}^{*})=1$ over each of the binary classifications $i$ at time $t$. The classification lines reflect criteria based on which decisions will be made. Imagine for example that $x_{1}$ corresponds to food and $x_{2}$ to water reward. Depending on the agent’s internal state, one takes precedence over the other, and the degree of preference is reflected in the slope of the line.
Figure 1: Disentangled representations and a framework to learn them. (a) A disentangled representation directly lends itself to OOD generalization: a downstream linear decoder that can differentiate ripe from unripe bananas can readily generalize to mangos, even though it has never been trained on mangos. (b) Overview of our multi-task classification framework. A ground truth $\mathbf{x}^{*}$ is sampled and Gaussian noise is added to arrive at observations $\{\mathbf{X}(1),...,\mathbf{X}(t)\}$. These observations are processed by the filter-based model illustrated graphically in Figure S5, maintaining a latent state $\mathbf{Z}(t)$. The latent state $\mathbf{Z}(t)$ is then used to produce classification outputs $\hat{Y}_{1}(t)$, $\hat{Y}_{2}(t)$. Theorem B.6 proves that $\mathbf{Z}(t)$ must encode the optimal estimator of $\mathbf{x}^{*}$ given the noisy observations, $\mu(t)$.
Criterion for Disentangled Representation Learning:
We investigate how solving the multi-task classification problem (Figure 1b) leads to agents learning disentangled representations of the latent ground truth $\mathbf{x}^{*}$ in its internal state $\mathbf{Z}(t)$. Specifically, we ask whether there exists a linear-affine transformation $(\mathbf{A},\mathbf{b})$ such that $\mathbf{x}^{*}=\mathbf{A}\mathbf{Z}(t)+\mathbf{b}$. Such a mapping would imply $\mathbf{Z}(t)$ linearly represents $\mathbf{x}^{*}$. If the rows of $\mathbf{A}$ are approximately orthogonal, the representation is disentangled.
@@ 3 Theoretical results
Here we provide conditions and guarantees for the emergence of disentangled representations in optimal multi-task classifiers with latent state $\mathbf{Z}(t)$ in the paradigm described in Section 2 and Figure 1. By “optimal multi-task classifier”, we refer to any agent or system whose outputs $\hat{\mathbf{Y}}(t)$ correspond to the correct posterior classification probabilities given the noisy, non-linearly transformed observations; that is, for each task $i=1,\dots,N_{\textit{task}}$
$\hat{Y}_{i}(t)$ $=\Pr\left(y_{i}(\mathbf{x}^{*})=1\mid f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))\right)$ (2)
$\hat{\mathbf{Y}}(t)$ $=g(\mathbf{Z}(t))$ (3)
The notion of optimality allows us to make precise statements about the informational content of the agent’s internal state $\mathbf{Z}(t)$. Let $\mathbf{C}\in\mathbb{R}^{N_{\textit{task}}\times D}$ be a matrix where each row is a decision boundary normal vector. Then
@@
If $\mathbf{C}\in\mathbb{R}^{N_{\textit{task}}\times D}$ is a full-rank matrix and $N_{\textit{task}}\geq D$ and noise $\sigma>0$, then
@@
    any optimal estimator of $\mathbf{y}(\mathbf{x}^{*})$ must encode a finite-sample, maximum likelihood estimate $\mu(t)$ of the ground truth evidence variable $\mathbf{x}^{*}$ in its latent state $\mathbf{Z}(t)$, and
@@
    if the activation function is sigmoid-like, $\mu(t)$ will be linearly decodable from $\mathbf{Z}(t)$, thus implying that $\mathbf{Z}(t)$ contains an abstract representation of $\mu(t)$ [59].
  3. 3.
    Furthermore, the representation is guaranteed to be disentangled (orthogonal) as $N_{\textit{task}}\gg D$ for random decision boundaries.
Specifically, $\mu(t)$ is the maximum likelihood estimate (MLE) of $\mathbf{x}^{*}$ given observations $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$. A closed-form expression for extracting $\mu(t)$ from $\mathbf{Z}(t)$ is given by:
@@
where $\Phi$ is the CDF of the normal distribution, $\sigma$ is the noise magnitude and $t$ the trial duration. Furthermore, if the activation function $g$ is of the sigmoid family of functions ($\tanh,\textup{sigmoid}$), then the term $\Phi^{-1}\big{(}g(\cdot))$ approximately cancels out, leading to:
$\mu(t)$ $\approx\underbrace{\frac{a_{g}\,\sigma}{\sqrt{t}}(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{Z}(t)}_{\text{Linear Function of$\mathbf{Z}(t)$}}+\underbrace{(\mathbf{C}^{\top}\mathbf{C})^{-1}\mathbf{C}^{\top}\mathbf{b}}_{\text{Affine Term}}$ (5)
where we have approximated $a_{\tanh}=\frac{2\sqrt{3}}{\pi}$ for $g=\tanh$ and $a_{\sigma}=0.5886$ for $g=\textup{sigmoid}$. For Gaussian IID noise, $\mu(t)$ is the sample mean of $\{\mathbf{X}(t)\}_{t=1}^{T}$, i.e. with non-linearity $f$ removed.
@@
Point 1 and Equation 4 are proven in Appendix B in Theorem B.6. Point 2 and Equation 5 are proven in Corollary B.7 for $\tanh$ and Corollary B.8 for sigmoid. Point 3 is proven in Corollary B.9. ∎
The key conceptual insight in the proof of Theorem 3.1 is that each of the multi-task classification probability estimates $\hat{Y}_{i}(t)$ represents an estimated projection distance between the MLE $\mu(t)$ and the given classification boundary $(\mathbf{c}_{i},b_{i})$. Once distances to classification boundaries are recovered, $\mu(t)$ can be inferred if the $N_{\text{task}}$ classification boundaries span the $D$-dimensional space of $\mathbf{x}^{*}$. Note that classification boundary distances are only guaranteed to be recoverable when there is non-zero noise $\sigma>0$ such that $\hat{Y}_{i}(t)$ does not saturate to 1 or 0, and thus still carries useful decision boundary distance information (see Lemma B.3). In fact, Equations 4 and 5 do not hold for $\sigma\to 0$, as they were derived by means of Bayesian estimation which assumes the presence of noise.
The linear approximation for decoding $\mu(t)$ from $\mathbf{Z}(t)$ in Equation 5 is enabled by the remarkable similarity between sigmoid functions and the Gaussian CDF $\Phi$. The sigmoid-like structure of $\Phi$ suggests many similar activations $g$ (e.g., softmax) would exhibit approximate linear decodability. For more exotic element-wise activation functions $g$, Equation 4 suggests $\mu(t)$ will be represented in $g(\mathbf{Z}(t))$ in the column space of $(\mathbf{C^{\top}C})^{-1}\mathbf{C}^{\top}$ potentially with some element-wise non-linearity.
Robustness to other noise distributions
Although we focus on Gaussian noise for our theoretical derivations, we note that a wide range of alternative noise distributions are permissible, including anisotropic Gaussian noise (Lemma B.10). The key property enabling us to recover distances from classification probabilities is the invertibility of the marginal posterior CDFs along the decision boundary normals. This property is not exclusive to Gaussian noise but is shared by a broader class of noise distributions, including but not limited to elliptical distributions (e.g., multivariate t-distributions, Laplace distributions), exponential power distributions, and other symmetric and unimodal distributions with invertible marginals. Therefore, the proof of Theorem 3.1 generalizes to any noise distribution satisfying the conditions outlined above (see Appendix B.4 for conditions and derivations). In fact, our experiments show robustness to anisotropic and autocorrelated noise, and mismatch between noise CDF and activation function (Section A.8).
4 Methods
We trained autoregressive models (RNNs, LSTMs, GPT-2 transformers) with latent state $\mathbf{z}(t)$, to output multi-task classifications $\mathbf{y}(\mathbf{x}^{*})\in\{0,1\}^{N_{\textit{task}}}$ given noisy and non-linearly mapped inputs $f(\mathbf{X}(1)),\dots,f(\mathbf{X}(t))$ (Figure 2). We subsequently trained linear probes $\mathbf{A}$ on $\mathbf{z}(t)$ to estimate $\mathbf{x}^{*}$, denoted $\hat{\mu}(t)=\mathbf{A}\,\mathbf{z}(t)$. We here focus on leaky RNNs, representing a brain area making decisions; for more details on GPT-2 experiments see Section A.5. The networks contain $N_{\textit{neu}}$ neurons, and their activations $\mathbf{z}(t)$ obey:
$\tau\dot{\mathbf{z}}=-\mathbf{z}+\left[\,\mathbf{W}_{\textit{rec}}\,\mathbf{z}+\mathbf{W}_{\textit{in}}\,\mathbf{x}_{\textit{in}}+\mathbf{b}\,\right]_{+}$ (6)
where $\mathbf{W}_{\textit{rec}}$ is the recurrent weight matrix, $\mathbf{W}_{\textit{in}}$ is the matrix carrying the input vector $\mathbf{x}_{\textit{in}}$, $\mathbf{b}$ is a unit-specific bias vector, $\tau$ is the neuronal time constant, $\left[.\right]_{+}$ is the ReLU applied element-wise and time dependencies have been dropped for brevity. We discretize Equation 6 using the forward Euler method for $T=20$ timesteps of duration $\Delta t=\tau=100\,\textrm{ms}$, which we find to be stable. The RNN’s output $\hat{\mathbf{y}}(t)\in\mathbb{R}^{N_{\textit{task}}}$ is given by $\hat{\mathbf{y}}(t)=g(\mathbf{W}_{\textit{out}}\,\mathbf{z}(t))$, where $\mathbf{W}_{\textit{out}}$ is a readout matrix and $g=\text{sigmoid}$ the output activation function applied elementwise.
Figure 2: Data generation and architecture. (a) The task is to simultaneously report whether the ground truth $\mathbf{x}^{*}$ lies above ($1$) or below (00) a number of classification lines. (b) Models (RNN depicted) are trained to report the outcome of all the binary classifications in a at the end of the trial (indicated by the fixation input turning 0).
The encoder $f$ is a 3-layer MLP with hidden dimensions $100,100,40$ and ReLU non-linearities, and it is randomly initialized and kept fixed during training as it represents a static mapping from latents to observations (observation map). An additional fixation input is directly passed to the hidden layer. It is 1 during the trial and turns 0 at the end of the trial, indicating that the network should report its decisions (Figure 2b). The fixation input precludes the RNN from learning a specific timing in its response. We refer to this kind of tasks as fixed reaction-time (RT). The network is trained with a cross-entropy loss and Adam default settings, except learning rate $\eta_{0}=10^{-3}$, to produce the target outputs $\mathbf{y}(\mathbf{x}^{*})$. By minimizing loss across trials, the network is incentivized to estimate $\hat{\mathbf{Y}}(t)=\Pr\{y_{i}(\mathbf{x}^{*})=1\}$. Table S1 summarizes all hyperparameters and their values, which are shared across all architectures.